# VoiceCloak: A Multi-Dimensional Defense Framework against Unauthorized Diffusion-based Voice Cloning

### v2

@@ 5.3.1. Effectiveness of Adversarial Identity Obfuscation
- - 1.99 45.00% 0.31 3.09 20.20% 3.38 1.35 34.82
$\checkmark$ - 2.42 31.80% 0.29 2.68 41.20% 3.30 1.30 34.03
- $\checkmark$ 2.23 23.00% 0.27 2.44 60.60% 3.28 1.29 33.83
- w/o Sem-free 2.28 26.36% 0.29 3.30 26.80% 3.00 1.36 33.28
$\checkmark$ $\checkmark$ 2.22 23.60% 0.27 2.10 57.80% 3.29 1.29 33.91
We first investigate the individual contributions of the loss components designed for Adversarial Identity Obfuscation (Section 4.2), namely $\mathcal{L}_{ID}$ and $\mathcal{L}_{ctx}$. The results are presented in Table 2. Employing $\mathcal{L}_{ID}$ markedly reduces the speaker verification accept rate. This confirms that directly manipulating the representation learning embedding space effectively disrupts recognizable identity features. Furthermore, the higher ASV (19.92%) when removing opposite-gender guidance indicates our psychoacoustically-motivated strategy provides directional control to specifically disrupt structural identity cues. The context divergence loss $\mathcal{L}_{ctx}$ also significantly contributes to defense, lowering both ASV and NISQA. This suggests that $\mathcal{L}_{ctx}$ interferes with the attention mechanism’s conditioning on the reference speaker’s style, leading to a boost in the overall DSR.
5.3.2. Effectiveness of Perceptual Fidelity Degradation
Table 3 shows the ablation study evaluating components designed for perceptual quality degradation. We observe that applying $\mathcal{L}_{score}$ alone forces the denoising trajectory away from the high-fidelity manifold via its amplified drift term. Consequently, NISQA drops to 2.68 and DTW distance increases. The semantic corruption loss $\mathcal{L}_{sem}$ demonstrates an even stronger individual impact on perceived quality. Its effectiveness stems from directly corrupting internal U-Net features, thereby impairing the model’s reconstruction of coherent, natural-sounding speech details. Critically, the necessity of guiding semantic features towards an incoherent state is evident from the ”w/o Sem-free” variant. Removing this noise-guidance component results in audio quality comparable to the baseline.
5.3.3. Perturbation Budget
Table 4. Performance of VoiceCloak under different perturbation budgets. ”Undefended” denotes voice conversion without VoiceCloak.
Budgets ($\epsilon$) Defense Effectiveness Imperceptibility
@@ 5.3.3. Perturbation Budget
Undefended - 76.49% - 3.96 - - - -
0.0005 2.01 22.00% 0.29 2.86 41.20% 4.25 0.48 45.03
0.0010 2.05 16.60% 0.28 2.65 57.00% 3.81 0.79 39.26
0.0020 2.12 11.40% 0.27 2.36 71.20% 3.21 1.29 33.53
0.0050 2.23 6.20% 0.25 1.95 84.00% 2.34 2.34 25.94
0.0100 2.32 1.20% 0.24 1.75 90.80% 1.74 3.40 20.18
We investigated the influence of the perturbation budget $\epsilon$ on both defense effectiveness and adversarial imperceptibility, with results summarized in Table 4. The findings reveal a clear trade-off: increasing e directly enhances defense effectiveness but conversely degrades audio imperceptibility. The excessive distortion compromises the usability of $x_{adv}$ for legitimate purposes. We determined $\epsilon=0.0020$ to be a suitable default budget, providing strong protection while preserving acceptable adversarial quality.
5.3.4. Impact of the Sampling Step
Table 5. Ablation study on the effect of different diffusion inference steps $T$. ”Avg. RT” indicates the average runtime per generated sample.
Infer. Steps $T$ Defense Effectiveness Avg. RT (s)
[…]
6 1.72 11.80% 1.83 84.20% 0.32
18 1.94 10.40% 2.08 81.80% 0.71
30 2.03 11.40% 2.13 78.80% 1.09
100 2.13 11.00% 2.37 71.80% 3.28
200 2.15 13.40% 2.37 69.20% 6.57
Table 5 investigates the impact of varying inference steps $T$, addressing the realistic scenario where an attacker’s generation timestep is unknown. The results demonstrate strong stability in defense effectiveness regardless of the timesteps. Notably, the DTW score, tends to increase with more inference steps. We attribute this to the cumulative amplification effect discussed in Section 4.3, whereby induced trajectory deviations compound over longer inference paths.
6. Conclusion
In this paper, we introduced VoiceCloak, a novel, multi-dimensional proactive defense framework designed to safeguard against unauthorized diffusion-based voice cloning. We established that a comprehensive defense should achieve the dual objectives of speaker identity obfuscation and perceptual quality degradation by exploiting targeted intrinsic vulnerabilities. To achieve identity obfuscation, it directly manipulates speaker representations and interferes with the attention mechanism’s understanding of contextual information via attention context divergence. For perceptual quality degradation, it employs Score Magnitude Amplification to steer the reverse denoising trajectory and utilizes noise-guided semantic corruption to promote semantic incoherence by disrupting internal feature processing. Experiments demonstrate the effective defense performance of VoiceCloak, thereby mitigating the potential harms associated with high-quality diffusion-based voice mimicry.
Acknowledgements.
This work is supported by the National Natural Science Foundation of China (No.62441237, No.62261160653, No.U2001202, No.62072480), the Guangdong Provincial Key Laboratory of Information Security Technology (No. 2023B1212060026).

### v4

@@ 5.3 Ablation Study
- - 1.99 45.00% 0.31 3.09 20.20%
$\checkmark$ - 2.42 31.80% 0.29 2.68 41.20%
- $\checkmark$ 2.23 23.00% 0.27 2.44 60.60%
- w/o Sem-free 2.28 26.36% 0.29 3.30 26.80%
$\checkmark$ $\checkmark$ 2.22 23.60% 0.27 2.10 57.80%
Table 3: Ablation study for Perceptual Fidelity Degradation. Checkmark ($\checkmark$) indicates the setting is used,”w/o” denotes the exclusion of the specified loss term.
We ablate the loss designed for Adversarial Identity Obfuscation, $\mathcal{L}_{ID}$ and $\mathcal{L}_{ctx}$, in Table 2. The results show that $\mathcal{L}_{ID}$ alone effectively reduces the ASV acceptance rate. This confirms that directly manipulating the representation learning embedding space effectively disrupts recognizable identity features. Removing the opposite-gender guidance from $\mathcal{L}_{ID}$ worsens ASV, confirming our psychoacoustically-motivated strategy provides effective direction for identity disruption. Separately, the context divergence loss also contributes significantly, lowering both ASV and NISQA by interfering with the attention mechanism’s condition injection.
Effectiveness of Perceptual Fidelity Degradation
Table 3 ablates the perceptual quality degradation losses. The results show that $\mathcal{L}_{score}$ alone is effective, degrading quality (lower NISQA, higher DTW) by forcing the denoising trajectory away from high-fidelity regions. The semantic corruption loss, $\mathcal{L}_{score}$, demonstrates an even stronger individual impact by directly corrupting internal U-Net features, thereby impairing the model’s reconstruction of coherent, natural-sounding speech details. Critically, the necessity of guiding semantic features towards an incoherent state is evident from the ”w/o Sem-free” variant.
5.4 Transferability
Target Models Defense Effectiveness
@@ 5.4 Transferability
DiffVC 2.12 11.40% 0.27 2.36 71.40%
DDDM-VC 1.67 16.80% 0.36 2.79 54.89%
DuTa-VC 2.41 13.77% 0.27 2.14 73.92%
Table 4: Transferability of the proposed defense method against unseen target models.
To demonstrate applicability, we extended our experiments to include two additional open-source diffusion-based VC models: DuTa-VC (Wang et al. 2023) and DDDM-VC (Choi, Lee, and Lee 2024). As shown in Table 4, VoiceCloak demonstrate favorable transferability to different models, achieving an average DSR of 66.7%. We attribute this transferability to: Targeting Common Vulnerabilities, as our method exploits fundamental mechanisms (e.g. attention, score prediction) that are shared across diffusion VCs.
5.5 Robustness
Figure 6: Resilience of VoiceCloak under more advanced robust scenario.
We evaluated the robustness of our method against four common distortions, with results presented in Figure 6. The results show consistent defense effectiveness, indicating resilience to real-world transformations and significantly outperforming the undefended baseline. Further details and analysis are available in Appendix.
5.6 Efficiency Discussion
Time (s) Mem. (GiB)
Ours 148.66 8.97
Table 5: Computational overhead analysis. ”Time”: average time to generate one sample. ”Mem.”: peak GPU usage.
As shown in Table 5, the protection process is a one-time, offline operation, and the required GPU memory is within the capacity of consumer-grade hardware, making VoiceCloak a practical defense method.
6 Conclusion
This paper introduced VoiceCloak, a comprehensive defense against unauthorized diffusion-based voice cloning (VC). Our framework achieves superior defense effectiveness by exploiting targeted intrinsic vulnerabilities within the diffusion process. Through strategies designed to disrupt attention-based conditional guidance, steer the denoising trajectory, and corrupt internal semantic representations, VoiceCloak effectively undermines the synthesis process. Extensive experiments validate its efficacy, demonstrating success in simultaneously obfuscating speaker identity and degrading audio quality to mitigate the threats of voice mimicry Extensive experiments validate its ability to significantly hinder voice cloning by simultaneously disrupting identity and degrading audio quality, thereby mitigating the threats of high-quality voice mimicry